# Generative AI for Intent-Driven Network Management in 6G: A Case Study on Hierarchical Learning Approach

### v1

@@ III Three-step Methodology for Intent-driven Network Management
$\bullet$ Intent execution: Intent execution is the final stage of IDN management, where validated intents are transformed into real-time network actions. This phase involves invoking appropriate network optimization policies to dynamically configure resources, and ensure that the desired service-level objectives are met efficiently. Intent execution can be realized through network function orchestration, AI-driven decision-making, or Reinforcement Learning (RL)-based adaptation. There are multiple ways to execute intents depending on the network architecture and management framework. The system can trigger predefined network applications such as traffic steering, beamforming, power control, and slicing management for intent execution. Also, LLMs can generate and refine network policies dynamically. Another approach for multi-operator environments can be to utilize RL-based agents that can negotiate optimal execution strategies to ensure efficient network optimization.
To represent this three-step methodology, we presented Fig.1 in the Introduction, which illustrates the step-by-step process of intent-driven network management. The process begins with intent processing, where network intents can be expressed in natural language, structured formats (JSON, XML), or multi-modal data such as topology diagrams and configuration files. To process the intents, different methodologies can be applied, including LLMs for natural language understanding, rule-based approaches using predefined logic, graph-based methods for structured representation of dependencies, or finite state machines to model intent workflows as a sequence of state transitions.
Once an intent is processed, it moves to intent validation which can be carried out using rule-based validation (checking against static policies), constraint-based validation (ensuring resource availability and compliance), predictive validation (using AI/ML models to anticipate network conditions), or RLHF to refine validation mechanisms over time.
After successful validation, the intent execution phase applies the validated intent to the network infrastructure. The execution can be performed using API-based deployment for direct implementation, Open RAN network applications for network-specific functionalities, or negotiation-based execution via multiple LLM agents to resolve potential conflicts between different intents before deployment.
@@
III-A Hierarchical RAN Integration of the Proposed IDN Management Architecture
@@
Modern hierarchical RAN architectures, such as Open RAN, use a multi-layered and modular design to support flexibility, scalability, and interoperability in next-generation wireless networks. These systems break down the traditional monolithic RAN into software-driven components that enable automation and AI-based optimization. In this work, we adopt a two-level architecture composed of a strategic controller and a tactical controller. The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations including handovers, resource allocation, and traffic steering. This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels, ensuring accurate interpretation and efficient execution of network intents.
@@
We provide an example of implementing the three-step IDN management workflow using a custom hierarchical RAN architecture in Fig. 2. A human operator can observe the current network status or query for any relevant network information before providing an intent. The LLM module in the picture is given as an example. The operator may use logs, or another query-based database to gather information instead of using an LLM. After that, the operator can provide an intent that is received by the user interface application in the top-level Intelligent RAN Controller (IRC) named strategic controller. It then passes through the intent validation application. This application can be designed to perform intent validation via rule-based static methodologies or GenAI-based predictive analysis. Lastly, at the bottom, there is a tactical controller responsible for intent execution. In Fig. 2, we provide an example of intent execution via network applications in the tactical controller. Based on the intent, multiple network applications such as traffic steering, power allocation, and cell sleeping can be initiated and orchestrated to fulfill an intent.
IV A Case Study with Hierarchical Learning for IDN Management
@@
Hierarchical application of GenAI algorithms can optimize complex decision-making by structuring learning at multiple levels. For instance, LLMs can interpret human intents, transformer-based predictors can validate network conditions, and decision transformers can execute intents. At the highest level, an LLM extracts key metrics and constraints from operator objectives. These insights guide a mid-tier forecasting model that predicts network behavior using historical and real-time data. At the lowest level, control algorithms generate policies for intent execution. This layered hierarchical approach improves efficiency, reduces computational overhead, and enables adaptive, real-time decisions
@@
Fig. 3 presents an end-to-end GenAI framework for IDN management, which forms the foundation of our case study. Data collection follows a hierarchical reinforcement learning approach with decision-making split across two levels. The higher-level controller defines goals based on operator intents, such as increasing throughput or reducing power usage. The lower-level controller selects appropriate network applications depending on real-time conditions like traffic load, energy consumption, and packet loss. Five applications are used in this study: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover management. Each of them is developed using DRL. Traffic steering application distributes traffic intelligently across Radio Access Technologys (RATs) to maintain QoS, optimizing both delay and throughput. Cell sleeping reduces power consumption by deactivating underutilized base stations based on traffic load and queue length. Beamforming calculates optimal steering angles and selects matching array vectors using user location and the power allocation application improves total throughput through smart resource distribution. Lastly, the handover management application enhances energy efficiency by learning adaptive handover decisions. Once selected, each application independently adjusts its parameters using its own learning policy. More details can be found in [4].
Once an action is taken, the system evaluates its effectiveness by assessing whether the applied changes contribute to the desired network improvements. This evaluation is based on a reward system that provides feedback on whether the selected applications led to better performance. Over time, the system collects a large number of these decision-making instances to form a dataset that contains detailed records of past actions, network conditions, and resulting outcomes. The dataset is obtained via a custom-built simulation environment that mimics real-world network scenarios. Over time, this dataset can be extended and updated with real-world traces and trajectories to enhance its realism and applicability.
@@
The collected data is structured into three distinct datasets. The first dataset is designed for fine-tuning an LLM and consists of query-response pairs and intent-based prompts to enable the model to accurately interpret operator-defined intents and provide relevant network insights. The second dataset comprises time series data, which includes key network parameters such as traffic load, power consumption, and packet loss percentage. This dataset is crucial for predictive intent validation, allowing the system to anticipate future network conditions and assess whether an intent can be successfully executed without degrading performance. The third dataset is a trajectory dataset used in the final stage, where control algorithms learn to optimize network application selection and orchestration. This dataset captures past decisions, network states, and corresponding outcomes to form the basis for training decision models that can make informed, goal-driven choices to achieve operator-defined objectives efficiently.
In this case study, we employ QLoRA to fine-tune an LLM (Llama 3.2 by Meta) to minimize computational resource requirements. Instead of traditional fine-tuning approaches that require high memory and processing power, QLoRA significantly reduces memory consumption without sacrificing performance. This fine-tuned model is further integrated with a RAG module  [^1] The RAG architecture is hybrid: it combines retrieval-based document fetching with the generative capabilities of the fine-tuned LLM., allowing it to access up-to-date network information and respond to dynamic queries with greater accuracy.
In the intent validation stage, the system evaluates whether the operator’s intent can be safely and effectively executed under current and predicted network conditions. This is done using a data-driven and predictive framework. First, the system continuously monitors key performance indicators including traffic load, packet loss percentage, and power consumption. These metrics are forecasted into future time slots using a transformer-based time series model that captures trends and relationships in temporal data. Once an intent is submitted, such as a request to improve throughput or reduce power usage, the system validates it against predicted network conditions. A multi-level check is performed where each intent is compared to predefined thresholds for the three metrics. For example, if a high throughput increase is requested during a period of already high traffic, the system may flag the intent as infeasible. A lookup table is used to cross-reference predicted values and QoS outcomes from past executions to enable fast and intelligent validation. This process helps ensure that only those intents are executed which are likely to maintain or improve overall network performance.
@@
Finally, the last part of this case study includes the Mamba architecture for intent execution via network application orchestration. In particular, we use decision Mamba proposed in [13]. We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates decision Mamba as the high-level decision mechanism and a Decision Transformer (DT) as the low-level control transformer.
@@
At a high level, decision Mamba serves as the goal-aware memory mechanism, responsible for retaining and retrieving critical past actions that have significantly contributed to achieving network optimization objectives. Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a Selective-State Space Modeling (SSSM) with a learnable dynamic memory retention mechanism. This allows the system to selectively remember only the most relevant past actions while discarding irrelevant ones. As a result, computational overhead gets reduced and decision efficiency improves.
@@
At the low level, the decision transformer [14] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the decision Mamba. Once Mamba identifies the most impactful past action, it is passed to the control transformer, which uses a sequence modeling approach to predict the optimal action for the current state. By conditioning decisions on goal awareness rather than predefined reward sequences, the control transformer ensures that each selected action aligns with the operator’s high-level intent.
@@
In this case study, when the operator provides an intent such as “Increase throughput by 10%”, the system sets this increased demand for throughput as its goal. It then recalls a past scenario with similar network conditions, such as high traffic load and moderate packet loss, where a specific combination of applications improved throughput. If the current network state does not support the intent, the system activates a feedback loop that either prompts the operator to revise the intent or uses the LLM to suggest feasible alternatives. The top-level GenAI block (decision Mamba) reviews the goal and the recent network context, retrieves the successful past action, and uses it to guide the present decision.
The control transformer described before has the current state of the network and the past action suggested by the top-level decision Mamba block. It uses this context to decide what to do next. It may decide to take the action of enabling a traffic steering application along with a power control application. This decision is tailored to current needs while being inspired by past success, aiming to fulfill the operator’s intent efficiently.
@@
To evaluate the effectiveness of HDMGA, we compare it against two baseline approaches: Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation. The HDTGA framework, which is our prior method, utilizes a hierarchical decision transformer architecture. A meta-transformer searches past trajectories to retrieve the most relevant past action, which is then used to guide the control transformer in predicting future actions. The HRL with intent validation baseline follows a conventional HRL approach with a dedicated intent validation mechanism [4].
@@
The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-RAT environment, serving $60$ users. The 5G New Radio (NR) operates at $3.5$ GHz (mid-band) and $30$ GHz (high-band) with bandwidths of $50$ MHz and $100$ MHz, respectively. The max transmission power of the 5G NR BS is $43$ dBm. On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $40$ MHz bandwidth and $38$ dBm maximum transmission power. Traffic types include video, gaming, voice, and vehicle-to-base station data traffic (Ultra-Reliable Low-Latency Communication (URLLC) use case), with packet inter-arrival times of $12.5$ ms, $40$ ms, $20$ ms, and $0.5$ ms, respectively, following Pareto, Uniform, and Poisson distributions.
In Fig. 4, we present spider plots of goal deviations for the proposed HDMGA and baseline methods. Delay-related goals (Fig. 4) include 2, 5, 7, and 11 ms. HDMGA consistently shows the lowest deviation, particularly excelling under stricter constraints. As delay targets relax (e.g., from 7 ms to 11 ms), all methods perform better due to increased flexibility, but HDMGA maintains superior accuracy. Similar trends are seen for energy-efficiency and throughput goals. Compared to HDTGA, HDMGA avoids attention overhead, reducing redundant computations. Additionally, Mamba’s linear-time processing enables efficient handling of long sequences, boosting throughput and making HDMGA more suitable for latency-sensitive applications.
Figure 4: Assessing the GenAI-based proposed three-step methodology in terms of delay goal deviations.
@@
Fig. 5 presents a plot showing that the action inference time of the proposed method is lower than that of the baselines. The actions illustrated are not sequential decisions taken under identical network conditions. Instead, each action corresponds to an independent inference made under distinct network states at specific time points during the simulation. Since network conditions (e.g., traffic load, power consumption, packet loss) and operator intents vary over time, different methods naturally select varying combinations of applications based on context. This variability reflects each method’s adaptability to dynamic RAN environments. The figure demonstrates that HDMGA achieves faster decision-making, which is essential for near-real-time network automation. This efficiency is enabled by its memory-guided design, where decision Mamba selectively retains and retrieves only the most relevant past actions. This eliminates the need for exhaustive searches over large trajectory buffers. All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (10 ms to 1 s) [15].
A major challenge in systems with many network applications is the large and complex action space, which increases training time and hampers generalization. Our prior work addressed this using an attention-based HRL framework to filter irrelevant actions [4]. In this work, the proposed HDMGA further mitigates the issue through supervised learning approach on an offline-collected dataset, eliminating the need for online training during deployment.

### v2

@@ III Three-step Methodology for IDN Management and its RAN Integration with GenAI
$\bullet$ Intent execution: Intent execution is the final stage of IDN management, where validated intents are transformed into real-time network actions. This phase involves invoking appropriate network optimization policies to dynamically configure resources and ensure that the desired service-level objectives are met efficiently. Intent execution can be realized through network function orchestration, AI-driven decision-making, or RL-based adaptation.
@@
III-A Hierarchical RAN Integration with the three-step IDN Management Architecture
@@
Modern hierarchical RAN architectures, such as Open RAN, use a multi-layered and modular design to support flexibility, scalability, and interoperability in next-generation wireless networks. These systems break down the traditional monolithic RAN into software-driven components that enable automation and AI-based optimization. In this work, we adopt a two-level architecture composed of a strategic controller and a tactical controller. The strategic controller manages long-term objectives like service quality, energy efficiency, and traffic patterns, while the tactical controller focuses on real-time operations, including handovers, resource allocation, and traffic steering. This layered approach helps implement intent processing, validation, and execution by supporting structured decisions across different time scales and abstraction levels.
@@
We provide an example of implementing the three-step IDN management workflow using a custom hierarchical RAN architecture in Fig. 1. A human operator can observe the current network status or query for any relevant network information before providing an intent. The LLM module in the picture is given as an example. The operator may use logs or another query-based database to gather information instead of using an LLM. After that, the operator can provide an intent that is received by the user interface application in the top-level Intelligent RAN Controller (IRC) named strategic controller. It then passes through the intent validation application. This application can be designed to perform intent validation via rule-based static methodologies or GenAI-based predictive analysis. Lastly, at the bottom, there is a tactical controller responsible for intent execution. In Fig. 1, we provide an example of intent execution via network applications in the tactical controller. Based on the intent, multiple network applications such as traffic steering, power allocation, and cell sleeping can be initiated and orchestrated to fulfill an intent.
IV A Case Study on Mamba Model for IDN Management
@@
Hierarchical use of GenAI algorithms can simplify complex decision-making by splitting learning across multiple layers. At the top layer, LLMs interpret human intents and extract key metrics and constraints. These insights guide a mid-tier time-series predictor that analyzes historical and real-time data to validate network conditions. At the lowest layer, control algorithms translate validated intents into executable policies. This layered design improves efficiency, lowers computational overhead, and supports adaptive, real-time decision-making.
@@
Fig. 2 shows the end-to-end GenAI framework for IDN management used in our case study. Data collection follows an HRL structure with two decision-making levels. The higher-level controller sets goals based on operator intents, such as raising throughput or lowering power use. The lower-level controller chooses the appropriate network application based on traffic load, energy efficiency, or packet loss. Five applications are used: traffic steering, cell sleeping, beamforming, power allocation, and energy-efficient handover. Each is implemented using DRL. Traffic steering distributes traffic across different radio access technologies to improve delay and throughput. Cell sleeping reduces power by disabling underutilized base stations. Beamforming computes optimal angles using user location, while power allocation improves total throughput through smart resource assignment. The handover management application enhances energy efficiency through adaptive handover decisions. After selection, each application independently adjusts its parameters using its own policy. More details are available in [3].
After each action, the system evaluates its impact by checking whether it improves the targeted network metrics. A reward mechanism measures how effective the selected applications were in enhancing performance. Over time, this process produces a large dataset containing past actions, network states, and observed outcomes. The dataset is generated using a custom simulator designed to mimic realistic network conditions and can later be enriched with real-world traces to further improve accuracy.
@@
The collected data is organized into three datasets. The first supports LLM fine-tuning and contains query–response pairs and intent-driven prompts to help the model interpret operator goals. The second dataset contains time-series measurements of parameters such as traffic load, packet loss, and power consumption. This data enables predictive intent validation by forecasting whether an intent is safe to execute. The third dataset is a trajectory set that trains the control algorithms. It records past decisions, states, and outcomes, enabling decision models to learn effective application selection that meets operator objectives.
In this case study, we adopt Mamba-Parameter Efficient Fine-Tuning (PEFT) to fine-tune a pretrained Mamba-based LLM for intent-driven RAN management [21]. Instead of updating all model parameters, which would require high memory and computing, Mamba-PEFT modifies only selected components inside the Mamba block. These include the input and output projections, specific state-space transitions, or a small number of additional state dimensions. By limiting the number of trainable parameters, the method reduces memory usage and keeps the fine-tuning process efficient while preserving the model’s long-context reasoning ability. Since Mamba processes sequences in linear time, the overall adaptation becomes more scalable than transformer-based alternatives. The fine-tuned model is integrated with a RAG module, which provides access to current network information, configurations, and logs. With both selective parameter adaptation and retrieval-augmented knowledge, the resulting Mamba-PEFT model delivers accurate and context-aware responses to operator queries while remaining suitable for use in hierarchical RAN environments.
In the intent validation stage, the system evaluates whether an operator’s intent can be executed safely under both current and predicted network conditions. KPIs such as traffic load, packet loss, and power consumption are continuously monitored and forecasted using Mamba4Cast [20]. Mamba4Cast is a Mamba-based zero-shot time-series foundation model built on selective state-space representations. It combines causal convolution layers with lightweight Mamba-based encoder blocks to capture multi-scale temporal patterns and generates multi-horizon forecasts in a single pass. When an operator submits an intent to increase throughput or reduce power usage, the predicted KPI trajectories are checked against predefined feasibility thresholds. If the forecasts indicate a risk of performance degradation, such as requesting higher throughput during a period of high expected load, the system marks the intent as infeasible. A lookup table that maps predicted KPI patterns to known quality-of-service outcomes from past executions further accelerates the decision. Through this combination of Mamba-based time-series forecasting and historical pattern matching, the validator ensures that only intents suitable for future conditions proceed to execution.
@@
Finally, the last part of this case study includes the Mamba architecture for intent execution via network application orchestration. In particular, we use the Decision Mamba proposed in [16]. We propose Hierarchical Decision Mamba with Goal Awareness (HDMGA), a hybrid hierarchical decision-making framework that integrates Decision Mamba as the high-level decision mechanism and a DT as the low-level control transformer.
@@
At a high level, Decision Mamba serves as the goal-aware memory mechanism, responsible for retaining and retrieving critical past actions that have significantly contributed to achieving network optimization objectives. Unlike conventional self-attention mechanisms, which perform exhaustive searches over past trajectories, Mamba employs a selective SSM with a learnable dynamic memory retention mechanism. This allows the system to selectively remember only the most relevant past actions while discarding irrelevant ones. As a result, computational overhead gets reduced, and decision efficiency improves.
@@
At the low level, the DT [19] functions as the control transformer, refining real-time actions based on the goal and past knowledge retrieved by the Decision Mamba. Once Mamba identifies the most impactful past action, it is passed to the control transformer, which uses a sequence modeling approach to predict the optimal action for the current state. By conditioning decisions on goal awareness rather than predefined reward sequences, the control transformer ensures that each selected action aligns with the operator’s high-level intent.
The intent format, inputs, and outputs used across the pipeline are as follows. For intent processing, the input is either a natural-language query requesting contextual network information or an optimization intent specifying a KPI and a desired magnitude. The output is either an LLM-generated informational response or a structured representation identifying the target KPI and magnitude, respectively. For intent validation, the input is the parsed optimization intent together with predicted network KPIs generated by the forecasting module, and the output is a feasibility decision indicating whether the intent can be safely executed under current or predicted conditions. For intent execution, the input is the desired magnitude of improvement of a KPI, and the output is a set of executable network actions, including the selected applications and their control parameters, deployed in the RAN controllers.
@@
In this case study, when the operator provides an intent such as “Increase throughput by 10%”, the system sets this increased demand for throughput as its goal. It then recalls a past scenario with similar network conditions, such as high traffic load and moderate packet loss, where a specific combination of applications improved throughput. If the current network state does not support the intent, the system activates a feedback loop that either prompts the operator to revise the intent or uses the LLM to suggest feasible alternatives. The top-level GenAI block (Decision Mamba) reviews the goal and the recent network context, retrieves the successful past action, and uses it to guide the present decision. The control transformer described before has the current state of the network and the past action suggested by the top-level decision Mamba block. It uses this context to decide what to do next. It may decide to take the action of enabling a traffic steering application along with a power control application. This decision is tailored to current needs while being inspired by past success, aiming to fulfill the operator’s intent efficiently.
Figure 3: Performance comparison of the proposed method against the baselines in terms of goal deviation (network delay and energy efficiency).
Figure 4: Performance comparison of the proposed method against the baselines in terms of throughput and reliability (QoS drift).
@@
To evaluate the effectiveness of HDMGA, we compare it against four baseline approaches. First of all, we adopt the deterministic rule-based method, which performs threshold-driven application triggering without any learning capability. This is referred to as Non-ML baseline in our performance graphs. Furthermore, we consider another baseline named StARformer where a transformer model is merged into an RL agent that predicts actions directly from recent state–action–reward tuples [23]. We address this as Transformer-RL (TRL) in our performance graphs. The last two baselines are Hierarchical Decision Transformer with Goal Awareness (HDTGA) and HRL with intent validation. The HDTGA framework, which is our prior method, utilizes a hierarchical decision transformer architecture. A meta-transformer searches past trajectories to retrieve the most relevant past action, which is then used to guide the control transformer in predicting future actions. The HRL with intent validation baseline follows a conventional HRL approach with a dedicated intent validation mechanism [3].
@@
The simulation setup in this case study consists of a macro cell surrounded by densely deployed small cells in a multi-Radio Access Technology (RAT) environment, serving $60$ users. The 5G New Radio (NR) operates at $3.5$ GHz (mid-band) and $30$ GHz (high-band) with bandwidths of $50$ MHz and $100$ MHz, respectively. The max transmission power of the 5G NR BS is $43$ dBm. On the other hand, Long-Term Evolution (LTE) operates at $800$ MHz with a $20$ MHz bandwidth and $38$ dBm maximum transmission power.
The simulator models the behavior of the RAN using a 3GPP-aligned design that includes realistic traffic patterns, mobility, and Radio Frequency (RF) impairments. Four traffic types are generated each having its own inter-arrival time and distribution, which creates a mix of bursty, periodic, and low-latency traffic. Traffic types include video, gaming, voice, and vehicle-to-base station data traffic (Ultra-Reliable Low-Latency Communication (URLLC) use case), with packet inter-arrival times of $12.5$ ms, $40$ ms, $20$ ms, and $0.5$ ms, respectively, following Pareto, Uniform, and Poisson distributions. User mobility follows 3GPP system-level guidelines, where User Equipment (UE) positions are updated every Transmission Time Interval (TTI) and movement directly affects path loss, Doppler shift, beam directions, and handover events. The RF environment includes Clustered Delay Line (CDL)-based multipath fading, shadowing, thermal noise, beam misalignment, and multi-cell interference. These impairments evolve with mobility and produce realistic fluctuations in signal-to-interference-plus-noise ratio, throughput, and delay. Together, the traffic models, mobility behavior, and RF effects create a reliable and reproducible simulation environment for evaluating intent-driven RAN management.
Our proof-of-concept is implemented using a multi-RAT MATLAB simulation (5G, LTE and CDL fading) integrated with Python-based Decision Mamba and Mamba-LLM modules. The LLM is fine-tuned using a 1.11.1k-sample dataset [19] of dynamic network queries, semi-static operational knowledge, and structured intents, while a Mamba-based predictor validates intent feasibility by forecasting load, power, and packet-drop trends.
As presented in Fig. 3, it can be observed that across both delay and energy-efficiency targets, the proposed HDMGA achieves the smallest goal deviations among all compared methods. This improvement is due to HDMGA ’s selective state-space modeling and goal-conditioned hierarchical design, which allow the agent to retain only the most relevant historical transitions and choose application combinations that directly contribute to meeting the operator’s intent. HDTGA performs moderately well but still shows higher deviations because its transformer-based attention weighs all past states uniformly and introduces unnecessary computational overhead. TRL experiences larger deviations since it lacks hierarchy, sub-goal decomposition, and selective memory, making it less effective when coordinating multiple network applications. HRL shows even higher deviations because it must explore the entire combinatorial action space without attention-based filtering or intent validation, which often results in suboptimal application choices under dynamic traffic conditions. The Non-ML heuristic baseline records the largest deviations for both metrics because its static threshold-based rules cannot adapt to changing network states or operator intent magnitudes. While only delay and energy-efficiency figures are presented, the same performance trend is consistently observed for throughput goal deviations as well.
The QoS drift rate represents how much the network’s performance deviates from the level expected by the operator when an intent is executed. The percentage is computed by averaging the deviation over a monitoring window that spans several TTIs. Fig. 4 compares the drift rates of all methods across different traffic loads. The proposed HDMGA maintains the lowest drift rates at every load level, indicating that it consistently meets the operator’s expected performance targets with minimal deviation. This reliability comes from its selective state-space modeling and goal-aware hierarchical design, which allow the agent to prioritize only the most relevant past transitions and choose the correct combination of applications even when the network becomes highly loaded.
The throughput results in Fig. 4 follow the same overall trend observed in the drift analysis. As the traffic load increases, HDMGA consistently achieves the highest throughput among all methods, reflecting its ability to select the most effective combination of RAN applications while respecting operator intent. Its selective memory mechanism helps preserve the most relevant context for scheduling and resource allocation, which leads to more efficient utilization of network capacity.
@@
Fig. 5 presents a plot showing that the action inference time of the proposed method is lower than that of the baselines. The actions illustrated are not sequential decisions taken under identical network conditions. Instead, each action corresponds to an independent inference made under distinct network states at specific time points during the simulation. Since network conditions (e.g., traffic load, power consumption, packet loss) and operator intents vary over time, different methods naturally select varying combinations of applications based on context. This variability reflects each method’s adaptability to dynamic RAN environments. The figure demonstrates that HDMGA achieves faster decision-making, which is essential for near-real-time network automation. This efficiency is enabled by its memory-guided design, where Decision Mamba selectively retains and retrieves only the most relevant past actions. All the methods in the figure maintain inference times within the near-real-time latency bounds defined by O-RAN specifications (1010ms to $1$s).